### v1

# CPCL: Cross-Modal Prototypical Contrastive Learning for Weakly Supervised Text-based Person Re-Identification

Abstract
Weakly supervised text-based person re-identification (TPRe-ID) seeks to retrieve images of a target person using textual descriptions, without relying on identity annotations and is more challenging and practical. The primary challenge is the intra-class differences, encompassing intra-modal feature variations and cross-modal semantic gaps. Prior works have focused on instance-level samples and ignored prototypical features of each person which are intrinsic and invariant. Toward this, we propose a Cross-Modal Prototypical Contrastive Learning (CPCL) method. In practice, the CPCL introduces the CLIP model to weakly supervised TPRe-ID for the first time, mapping visual and textual instances into a shared latent space. Subsequently, the proposed Prototypical Multi-modal Memory (PMM) module captures associations between heterogeneous modalities of image-text pairs belonging to the same person through the Hybrid Cross-modal Matching (HCM) module in a many-to-many mapping fashion. Moreover, the Outlier Pseudo Label Mining (OPLM) module further distinguishes valuable outlier samples from each modality, enhancing the creation of more reliable clusters by mining implicit relationships between image-text pairs. Experimental results demonstrate that our proposed CPCL attains state-of-the-art performance on all three public datasets, with a significant improvement of 11.58%, 8.77% and 5.25% in Rank@1 accuracy on CUHK-PEDES, ICFG-PEDES and RSTPReid datasets, respectively. The code is available at https://github.com/codeGallery24/CPCL.
Figure 1: Illustration of instance-wise contrastive learning and cross-modal prototypical contrastive learning. (a): It attempts to directly draw the image and text pairs closer together. (b): Each visual and textual instance is assigned a prototype corresponding to modality, according to its pseudo label.
1 Introduction
Text-based Person Re-Identification (TPRe-ID) is designed to retrieve all images of a pedestrian across multiple cameras based on a textual descriptionLi et al. (2017b); Ding et al. (2021); Zhu et al. (2021). A widely adopted framework for this task involves projecting the image and text input to a common latent space through separate encoders and aligning the features of the same identity in the latent space. The identity annotation is crucial for mitigating intra-class variations in the alignment process. On one hand, the annotation helps mitigate intra-modal feature variations. For example, images of a person could be captured by different cameras with varying illumination, view, and background. On the other hand, identity labels narrow down the cross-modal semantic gaps among input images and texts. However, obtaining identity annotations for the dataset is expensive and time-consuming. In a more practical setting, weakly supervised TPRe-ID Zhao et al. (2021) is considered, where the training set only contains text-image pairs without any identity annotations.
The absence of identity labels consequently introduces intra-class variations at both cross-modal and intra-modal levels. To address this challenge, CMMT Zhao et al. (2021) employed a clustering refinement method to refine pseudo labels, and a Text-IoU method was utilized to create soft labels for cross-modal matching. CAIBCWang et al. (2022a) proposed an instance-level matching method to avoid over-reliance on color information to enhance the performance of weakly supervised TPRe-ID. Although these methods have led to great performance advances in weakly supervised TPRe-ID, they suffer from the following issues. Firstly, the instance-wise contrastive loss treats two samples as a negative pair as long as they are from different pseudo labels and leads to an embedding space where all instances are well-separated, and each instance is locally smoothLi et al. (2021). It ignores the prototypical features of each person, which substantially hindered their further improvements. Secondly, they employed pre-trained unimodal models, such as ResNet-50 and Bi-LSTM as image and text encoders, respectively. However, they lack the necessary underlying alignment capabilities due to training on unimodal datasets. Thirdly, the soft labels generated by the Text-IOU were based solely on the percentage of overlapping noun phrases in the text description, which is unreliable and does not encourage the model to learn the overall semantic information of the caption.
To address the aforementioned challenges, inspired by human cognitive processRoss and Makin (1999); Lakoff (2007) and progress in other visual taskSnell et al. (2017); Li et al. (2021), we propose a Cross-Modal Prototypical Contrastive Learning (CPCL) method to guide the mulit-modal leaning. Intuitively, humans have the metacognitive ability to access knowledge to classify a sample as a certain type and then use their knowledge of this type to determine how to deal with itFleming and Dolan (2012). We propose a novel learning scheme where the visual samples of each person are assigned a visual prototype (and similarly, textual samples are assigned a textual prototype) to leverage this cognitive behaviour mechanism. A prototype is defined as a representative embedding in which samples around a single prototype representation for each modal class. The prototype acts as a bridge, overcoming intra-modal feature variations while mitigating the cross-modal semantic gaps, as illustrated in Fig. 1(b).
Furthermore, we cluster image samples and text samples separately to get a better initialization of the prototypes and assign each sample a pseudo label by DBSCAN algorithmEster et al. (1996). Considering that the prototypical features are constantly changing as the model updates, we propose the Prototypical Multi-modal Memory (PMM) module, which dynamically maintains the prototypical features so that the prototypical features can be kept up-to-date. To mitigate intra-class variation, we propose a Hybrid-level Cross-modal Matching (HCM) module. HCM calculates the Prototypical Contrastive Matching (PCM) loss between instance features and prototypical features of different modalities provided by PMM, along with Instance-level Cross-modal Projection Matching (ICPM) loss in each mini-batch with positive and negative samples obtained according to the pseudo labels. During the training process, we discovered a large number of un-clustered samples. We proposed an Outlier Pseudo Label Mining (OPLM) module to identify valuable outlier samples from each modality, enhancing the clustering quality by mining implicit relationships between image-text pairs. To obtain a better-aligned latent space of visual and textual encoders, we leverage CLIPRadford et al. (2021) as the backbone for our framework.
Our main contributions can be summarized as follows:
  • We propose a Cross-Modal Prototypical Contrastive Learning (CPCL) method designed to mitigate the intra-modal feature variations and cross-modal semantic gaps. This learning scheme is inspired by human cognitive behaviour.
  • We propose a Prototypical Multi-modal Memory (PMM) for Hybrid-level Cross-modal Matching (HCM), which helps CLIP establish the associations between heterogeneous modalities of image-text pairs belonging to the same class.
  • We introduce an Outlier Pseudo Label Mining (OPLM) module that further distinguishes valuable outlier samples by leveraging the implicit relationships between image-text pairs.
  • We have conducted extensive experiments that demonstrate our proposed CPCL achieves state-of-the-art results on three public benchmark datasets with up to average 8.53% Rank@1 gains. To the best of our knowledge, our work marks the first introduction of the CLIP pre-trained model to weakly supervised TPRe-ID.
Figure 2: Our proposed Cross-modal Prototypical Contrastive Learning assigned each visual and textual sample a prototype corresponding to modality, according to its pseudo label. PMM: Prototypical Multi-modal Memory module dynamically maintains the prototypical features. HCM: Hybrid-level Cross-modal Matching module consists of Prototypical Contrastive Matching (PCM) loss and Instance-level Cross-modal Projection Matching (ICPM) loss. OPLM: Outlier Pseudo Label MIning module identifies valuable un-clustered samples by mining implicit relationships between image-text pairs.
2 Related Work
2.1 Text-based Person Re-Identification
Text-based Person Re-Identification (TPRe-ID) differs from the classic Re-ID problemZheng et al. (2015); Ge et al. (2020); Zhao et al. (2022) as its objective is to identify pedestrian images based on provided text description. TPRe-ID was originally proposed by LiLi et al. (2017b), which utilized a CNN-RNN network to learn cross-modal features at the global level. Early worksLi et al. (2017a, b); Chen et al. (2018) employed VGGSimonyan and Zisserman (2014) and LSTMHochreiter and Schmidhuber (1997) for learning visual-textual representations. Subsequently, these representations were aligned by using a matching loss function. Building on this, ResNetHe et al. (2016) and BERTKenton and Toutanova (2019) were adopted for feature extraction, along with novel cross-modal matching losses to align global image-text features in a joint embedding space. Fine-grained approachesWu et al. (2021); Chen et al. (2022); Wang et al. (2022a) consider local and global matching relations. With increasing attention to CLIPRadford et al. (2021), more and more effortsYan et al. (2022); Jiang and Ye (2023); Bai et al. (2023) have endeavored to leverage CLIP’s powerful knowledge to establish an effective mapping relationship between images and texts. However, the aforementioned methods require identity annotations that are costly and time-consuming to obtain. To address this limitation, CMMTZhao et al. (2021) introduces weakly supervised text-based person re-identification. CMMT proposed MPLR to mine valuable un-clustered samples while employing Text-IOU to guide the model in learning image-text matching and to mine challenging samples.
2.2 Contrastive Learning
Contrastive learningHadsell et al. (2006) is a self-supervised learning method that learns discriminative feature representations by learning the similarities and dissimilarities among samples, leading to remarkable achievements in various domainsYe et al. (2019); He et al. (2020); Chen et al. (2020); Chen and He (2021); Misra and van der Maaten (2020). These methods are based on instance-wise contrastive learning and need to sample a sufficiently large number of negative instances. Though PCLLi et al. (2021) proposed prototypical contrastive learning to encourage the representation to encode the semantic structure of data, it was evaluated on the unimodal dataset and did not explore the complex relation between different modal samples in the weakly supervised TPRe-ID task. CLIPRadford et al. (2021) aligns images and texts closely in a common embedding space through cross-modal contrastive learning loss. Weakly supervised TPRe-ID is a fine-grained task and CLIP does not perform well if it is directly utilized on this task.
3 Method
3.1 Overview
The core of the proposed CPCL is to bring instance-level samples closer to their prototypes and then pull two modal prototypes of the same person together. The complete pipeline of our proposed method is illustrated in Fig. 2. Initially, we input all image-text pairs into the image and text encoder respectively, and extract features by CLIPRadford et al. (2021). Secondly, we use the DBSCANEster et al. (1996) algorithm to cluster images and texts separately, and Outlier Pseudo label Mining (OPLM) will identify valuable outlier samples. Thirdly, initialize Prototypical Multi-modal Memory (PMM) and train modal by Hybrid-level Cross-modal Matching (HCM) loss. Last, the outlier samples that OPLM does not mine in step three will be used to further train the model.
3.2 Prototpcial Multi-modal Memory
Humans can be capable of classifying a sampler as a certain type and then use their knowledge of this type to determine how to deal with itFleming and Dolan (2012). The proposed cross-modal prototypical contrastive learning simulates this cognitive behavior mechanism by bringing instance-level samples closer to their corresponding prototypes. Considering that the prototypical features are constantly updating as the training process, we propose the Prototypical Multi-modal Memory (PMM) module to dynamically maintain the prototypical features and keep them up-to-date.
3.2.1 Memory Initialization
Firstly, we employ the Image encoder to extract all image features, resulting in $F^{v}=\{f^{v}_{1},\ldots,f^{v}_{n_{v}}\}$. The superscript $v$ indicates that it is a visual feature and $n_{v}$ denotes the number of images. Then we utilize a clustering algorithm to cluster all image features and obtain pseudo labels $Y^{v}=\{y^{v}_{1},\ldots,y^{v}_{n_{v}}\}$.
To initialize the prototypical features for each class among the visual samples, we calculate a prototypical feature with pseudo label $i$ based on
$c^{v}_{i}=\frac{1}{\|Y^{v}_{i}\|}\sum_{f^{v}\in Y^{v}_{i}}f^{v},$ (1)
where $c^{v}_{i}$ represents the cluster centroid feature with pseudo label $i$ in the image memory, while $Y^{v}_{i}$ denotes a set of all image samples with the pseudo label $i$. Furthermore, the symbol |||||$\cdot$||| | ⋅ | | indicates the number of elements in a set. The initialization process for text memory follows the same procedure as image memory.
In conclusion, we perform separate clustering for both images and texts, which yields respective corresponding pseudo labels for each modality.
3.2.2 Memory momentum updating
After each training iteration, both image and text sample features within the given mini-batch are utilized to momentum update the prototypical features stored in prototypical memory, following the pseudo labels derived from clustering. Specifically, the feature of a certain image sample within the mini-batch, denoted as $f^{v}_{i}$, we perform the update to the image memory using
$c^{v}_{i}\leftarrow m^{v}c^{v}_{i}+(1-m^{v})f^{v}_{i},$ (2)
where $m^{v}$ represents the image memory momentum update coefficient, which controls the momentum update rate of class center samples. Similarly, For each text feature $f^{t}_{i}$ within the mini-batch, the text memory will also be momentum updated following the protocol as mentioned above. In our paper, we set $m^{v}$ and $m^{t}$ as 0.9 as default in our experiments.
3.3 Hybrid-level Cross-modal Matching
3.3.1 Protopypical Contrastive Matching
We employ two optional Protopypical Contrastive Matching (PCM): cross-modal PCM and single-modal PCM, denoted as $PCM_{c}$ and $PCM_{s}$, respectively. The formulation of the $PCM_{c}$ loss for image-to-text direction is as follows.
$\mathcal{L}^{v}_{c}=-\log\frac{\exp\left(f^{v}\cdot c^{t+}/\tau^{v}\right)}{\sum_{i=1}^{N^{t}}\exp\left(f^{v}\cdot c^{t}_{i}/\tau^{v}\right)}$ (3)
where $f^{v}$ denotes the image feature. The parameters $\tau^{v}$ and $N^{t}$ represent the learnable temperature parameter for the image memory and the number of cluster categories for captions, respectively. Symmetrically, the $PCM_{c}$ loss for text-to-image direction, denoted by $\mathcal{L}^{t}_{c}$ is derived by interchanging the roles of $f^{v}$ with $f^{t}$ and $c^{t}$ with $c^{v}$ within the framework of equation Eq. 3.
In summary, the bi-directional $PCM_{c}$ loss is calculated by:
$\mathcal{L}_{pcm}^{c}=\mathcal{L}^{v}_{c}+\mathcal{L}^{t}_{c}$ (4)
Similarly, $PCM_{s}$ loss $\mathcal{L}^{s}_{pcm}$ is also bi-directional and composed of $\mathcal{L}^{v}_{s}$ and $\mathcal{L}^{t}_{s}$. For example, replacing the $c^{t+}$ in Eq. 3 with $c^{v+}$ creates $\mathcal{L}^{v}_{s}$. In our paper, cross-modal PCM is adopted by default, $\mathcal{L}_{pcm}=\mathcal{L}_{pcm}^{c}$.
3.3.2 Instance-level Cross-modal Project Matching
In addition to matching at the prototypical level, instance-level matching is also necessary. Inspired by the CMPMZhang and Lu (2018), we applied Instance-level Cross-modal Project Matching (ICPM) loss for alignment at the instance level. Specifically, we calculate the probability of cross-modal matching pairs by the following softmax function:
$p_{i,j}=\frac{\exp\left(f^{v}_{i}\cdot f^{t}_{j}/\tau\right)}{\sum_{k=1}^{N}\exp\left(f^{v}_{i}\cdot f^{t}_{k}/\tau\right)},$ (5)
where footnote $i$ and $j$ denote the indexes of features in a mini-batch. $N$ is the number of image-text pairs. The pseudo cross-modal matching probability is $q_{i,j}=y_{i,j}/\sum_{k=1}^{N}y_{i,k}$. $y_{i,j}=1$ means that $(f^{v}_{i},f^{t}_{j})$ is a matched pair from the same pseudo label, while $y_{i,j}=0$ indicates the unmatched pair. Then the ICPM loss is computed by
$\mathcal{L}_{i2t}^{icpm}=KL(\mathbf{p_{i}}\|\mathbf{q_{i}})=\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{N}p_{i,j}\log(\frac{p_{i,j}}{q_{i,j}+\epsilon}),$ (6)
where $\epsilon$ is a small number to avoid numerical problems. Symmetrically, the ICPM loss from text to image $\mathcal{L}_{t2i}^{icpm}$ can be formulated by exchanging $f^{v}$ and $f^{t}$ in Eqs. 5 and 6. The bi-directional ICPM loss is calculated by:
$\mathcal{L}_{icpm}=\mathcal{L}_{i2t}^{icpm}+\mathcal{L}_{t2i}^{icpm}.$ (7)
Overall Loss By combining the losses defined above, the final objective of CPCL is formulated as:
$\mathcal{L}_{\textrm{overall}}=\mathcal{L}_{pcm}+\mathcal{L}_{icpm}$ (8)
Figure 3: Illustration of Outlier Pseudo Label Mining (OPLM). 1) For each outlier instance a, we group all the paired samples as a set B and eliminate any outlier samples from B. 2) If B is not empty, we identify the nearest sample set to B within the same modality and denote it as C. 3) Next, we search the set of paired instances D of C in the other modality and remove any outlier samples from D. 4) If D is not empty, we assign the class of the samples closest to a as the class of a.
3.4 Outlier Pseudo Label Mining
We propose an Outlier Pseudo Label Mining (OPLM) module that leverages existing image-text pairs information to mine valuable outlier samples, as illustrated as Fig. 3. OPLM consists of two stages: the refined learning stage and the supplementary learning stage, respectively.
3.4.1 The Refined Learning Stage
At the refined learning stage, the OPLM algorithm needs to be applied separately to both visual and textual modalities to identify valuable outlier samples from each modality. Consequently, it can be divided into two distinct components: $OPLM_{v\rightarrow t}$ and $OPLM_{t\rightarrow v}$.
For example, in $OPLM_{v\rightarrow t}$, for an outlier image sample $v^{o}_{i}$, we can find the corresponding text description according to the known image text corresponding index. However, it should be noted that there may be more than one paired text for an image. Therefore, its paired captions are
$P^{t}=PIS(v^{o}_{i})\ ,$ (9)
where $P^{t}=\{t_{1},\ldots,t_{p}\}$ represents that there are $p$ pieces of caption paired with this image, and $PIS(\cdot)$ denotes the process of paired instance searching in another modality. Subsequently, we proceed to filter $P^{t}$ by eliminating the outlier samples and obtain $P^{t}_{\textrm{filter}}$. If $||P^{t}_{\textrm{filter}}||=0$, $v^{o}_{i}$ remains as a outlier sample. Conversely, if $P^{t}_{\textrm{filter}}=\{t_{1},\ldots,t_{k}\}$, it indicates that there are $k$ clustered samples. We then process all the samples in the collection $P^{t}_{\textrm{filter}}$ to identify their corresponding images:
$P^{v}=\{PIS(t_{i})\mid t_{i}\in P^{t}_{\textrm{filter}}\ ,i=1,\ldots,k\}\ ,$ (10)
where, $P^{v}=\{v_{1},\ldots,v_{k}\}$ denotes $k$ images corresponding to $P^{t}_{\textrm{filter}}$ in a one-to-one manner. Following this, $P^{v}$ undergoes a filtering process where outlier samples are removed, resulting in $P^{v}_{\textrm{filter}}$. If $||P^{v}_{\textrm{filter}}||=0$, then $v^{o}_{i}$ is kept as a outlier sample. However, if $P^{v}_{\textrm{filter}}=\{v_{1},\ldots,v_{q}\}$, it signifies the presence of $q$ clustered image samples. Subsequently, the distance from $v^{o}_{i}$ to all samples in $P^{v}_{\textrm{filter}}$ is calculated, and the nearest sample is identified as $v_{i}$
$v_{i}=\mathop{\arg\max}\limits_{v_{i}\in P^{v}_{\textrm{filter}},v_{i}\not\in U^{v}}\langle v^{o}_{i}\cdot v_{i}\rangle,i=1,\ldots,q\ ,$ (11)
[…]
$C^{v}_{k}\leftarrow\{C^{v}_{k},v^{o}_{i}\}\ ,$ (12)
[…]
If $v_{i}$ is still an outlier sample, it will be added to the collection $U^{v}$. The process will then return to formula Eq. 11 for re-selection until $P^{v}_{\textrm{filter}}=U^{v}$. If the outlier sample $v^{o}_{i}$ remains unassigned to a specific category, it will be discarded, and the algorithm will proceed to the next outlier sample. During the process of Outlier Pseudo Label Mining (OPLM), all the outlier samplers in both modalities are traversed.
3.4.2 The Supplementary Learning Stage
At the supplementary learning stage, the image-text pairs that are not mined by the refined learning stage compose the training dataset. The main objective loss function is ITC loss. For each image feature $f^{v}_{i}$, we obtain an InfoNCEOord et al. (2018) loss between its image feature $f^{v}_{i}$ and all text features $f^{t}_{j}$ in the batch,
$\mathcal{L}_{\textrm{i2t}}^{itc}=-\log\frac{\exp\left(f^{v}_{i}\cdot f^{t}_{i}/\tau\right)}{\sum_{k=1}^{N}\exp\left(f^{v}_{i}\cdot f^{t}_{k}/\tau\right)},$ (13)
where the $N$ denotes the batch size, and $\tau$ denotes the hyperparameter of temperature. Symmetrically, for each text feature, the InfoNCE loss is denoted as $\mathcal{L}_{\textrm{t2i}}^{itc}$. The total loss function for ITC loss is defined as:
$\mathcal{L}_{\textrm{itc}}=\mathcal{L}_{\textrm{i2t}}^{itc}+\mathcal{L}_{\textrm{t2i}}^{itc}.$ (14)
For convenience, we will refer to an OPLM that contains only the first stage as a One-stage OPLM and an OPLM that contains two stages as a Two-stage OPLM.
Method Type Ref Image Enc. Text Enc. Rank-1 Rank-5 Rank-10 mAP mINP
Fully Supervised Text-based Person Re-Identification
DSSLZhu et al. (2021) L MM21 RN50 BERT 59.98 80.41 87.56 - -
SSANDing et al. (2021) L arXiv21 RN50 LSTM 61.37 80.15 86.73 - -
LBULWang et al. (2022b) L MM22 RN50 BERT 64.04 82.66 87.22 - -
Han et al.Han et al. (2021) G BMVC21 CLIP-RN101 CLIP-Xformer 64.08 81.73 88.19 60.08 -
CAIBCWang et al. (2022a) L MM22 RN50 BERT 64.43 82.87 88.37 - -
AXM-NetFarooq et al. (2022) L MM22 RN50 BERT 64.44 80.52 86.77 58.73 -
IVTShu et al. (2022) G ECCVW22 ViT-Base BERT 65.59 83.11 89.21 - -
CFineYan et al. (2022) L arXiv22 CLIP-ViT BERT 69.57 85.93 91.15 - -
IRRAJiang and Ye (2023) G CVPR23 CLIP-ViT CLIP-Xformer 73.38 89.93 93.71 66.13 50.24
RaSaBai et al. (2023) G IJCAI23 ALBEF-ViT ALBEF-BERT 76.51 90.29 94.25 69.38 -
Weakly Supervised Text-based Person Re-Identification
CMMTZhao et al. (2021) G ICCV21 ResNet LSTM 57.10 78.14 85.23 - -
CMMT(CLIP-ViT-B/16)* G ICCV21 CLIP-ViT CLIP-Xformer 59.57 79.53 86.53 54.66 39.78
CAIBCWang et al. (2022a) G MM22 ResNet BERT 58.64 79.02 85.93 - -
Baseline (CLIP-ViT-B/16) G - CLIP-ViT CLIP-Xformer 58.45 78.87 85.30 54.14 39.83
CPCL (Ours) G - CLIP-ViT CLIP-Xformer 70.03 87.28 91.78 63.19 47.54
Table 1: Performance comparisons with SOTA methods on CUHK-PEDES. “G” and “L” correspond to the global-matching and local-matching methods, respectively. (*) the implementation is based on the CMMT’s paper.
4 Experiment
We conduct experiments on three TPRe-ID datasets: CUHK-PEDESLi et al. (2017b), ICFG-PEDESDing et al. (2021) and RSTPReidZhu et al. (2021). The introduction of each dataset and evaluation metrics are shown in the Appendix.
Method Type Rank-1 Rank-5 Rank-10 mAP mINP
Fully Supervised Text-based Person Re-Identification
Dual PathZheng et al. (2020) G 38.99 59.44 68.41 - -
ViTAAWang et al. (2020) L 50.98 68.79 75.78 - -
SSANDing et al. (2021) L 54.23 72.63 79.53 - -
IVTShu et al. (2022) G 56.04 73.60 80.22 - -
CFineYan et al. (2022) L 60.83 76.55 82.42 - -
IRRAJiang and Ye (2023) G 63.46 80.25 85.82 38.06 7.93
RaSaBai et al. (2023) G 65.28 80.40 85.12 41.29 -
Weakly Supervised Text-based Person Re-Identification
CMMT(CLIP-ViT-B/16)* G 54.27 71.17 77.86 33.17 5.73
Baseline (CLIP-ViT-B/16) G 53.83 73.33 80.46 30.60 5.35
CPCL (Ours) G 62.60 79.07 84.46 36.16 6.31
Table 2: Performance comparisons with SOTA methods on ICFG-PEDES dataset.
Method Type Rank-1 Rank-5 Rank-10 mAP mINP
Fully Supervised Text-based Person Re-Identification
DSSLZhu et al. (2021) G 39.05 62.60 73.95 - -
SSANDing et al. (2021) L 43.50 67.80 77.15 - -
LBULWang et al. (2022b) L 45.55 68.20 77.85 - -
IVTShu et al. (2022) G 46.70 70.00 78.80 - -
CFineYan et al. (2022) L 50.55 72.50 81.60 - -
IRRAJiang and Ye (2023) G 60.20 81.30 88.20 47.17 25.28
RaSaBai et al. (2023) G 66.90 86.50 91.35 52.31 -
Weakly Supervised Text-based Person Re-Identification
CMMT(CLIP-ViT-B/16)* G 52.25 76.45 84.55 41.98 22.00
Baseline (CLIP-ViT-B/16) G 53.10 75.70 83.60 38.65 21.06
CPCL (Ours) G 58.35 81.05 87.65 45.81 23.87
Table 3: Performance comparisons with SOTA methods on RSTPReid dataset.
No. Methods Components CUHK-PEDES
OPLM $\mathcal{L}_{pcm}$ $\mathcal{L}_{icpm}$ Rank-1 Rank-5 Rank-10
0 Baseline 58.45 78.87 85.30
1 +OPLM+$\mathcal{L}_{id}$ 61.44 81.48 88.56
2 +$\mathcal{L}_{pcm}$ 65.14 84.62 90.51
3 +$\mathcal{L}_{icpm}$ 65.08 83.38 89.30
4 +$\mathcal{L}_{pcm}$+$\mathcal{L}_{icpm}$ 68.76 86.76 91.23
5 +OPLM+$\mathcal{L}_{icpm}$ 67.99 86.08 91.63
6 +OPLM+$\mathcal{L}_{pcm}$ 68.24 86.06 91.26
7 CPCL 70.03 87.28 91.78
Table 4: Ablation studies on our proposed components of CPCL.
No. Methods Components CUHK-PEDES
$PCM_{c}$ Memory update Two-stage OPLM Rank-1 Rank-5 Rank-10
0 $PCM_{s}$ 68.24 86.06 91.23
1 no update 69.10 86.18 91.19
2 One-stage OPLM 68.76 86.11 91.26
3 CPCL 70.03 87.28 91.78
Table 5: Ablation studies on optional components of CPCL.
4.1 Implementation Details
We adopt a pretrained CLIP-ViT-B/16Radford et al. (2021) as the backbone. The batch size is set to 128, and the model is trained with Adam optimizerKingma and Ba (2015) for 60 epochs with a learning rate initialized to $1\times 10^{-5}$ and cosine learning rate decay. In the beginning, the 5 warm-up epochs linearly increase the learning rate from $1\times 10^{-6}$ to $1\times 10^{-5}$. Following the clustering-based method CMMTZhao et al. (2021), we use DBSCANEster et al. (1996) for clustering before each epoch. Note that we utilize the CLIP-ViT-B/16 model fine-tuned with ITC loss as the Baseline. More details can be found in Appendix.
4.2 Comparison with State-of-the-art Methods
As CMMTZhao et al. (2021) is only tested on the CUHK-PEDES dataset and does not public its training code, to conduct the comparison of the backbone and additional module between CMMT and our method, we implemented CMMT with CLIP as the backbone to access its performance on ICFG-PEDES and RSTPeid.
4.2.1 Comparison with Weakly supervised TPRe-ID Methods
We compared our method with prior methods on three datasets, as shown in Table 1, Table 2 and Table 3, respectively. Our method CPCL achieves the SOTA performance in terms of all metrics, outperforming existing methods by a large margin. Specifically, compared with the Baseline, CPCL gains a significant Rank@1 improvement of 11.58%, 8.77% and 5.25% on the three datasets, respectively.
4.2.2 Comparison with Fully Supervised TPRe-ID Methods
To see the gaps between weakly and fully supervised TPRe-ID, we also list the performance of the fully supervised TPRe-ID methods in the 1st group on three datasets, as shown in Table 1, Table 2 and Table 3 respectively. Compared with current RaSaJiang and Ye (2023), CPCL is lower than 3.35%, 0.86% and 1.85% in Rank@1 accuracy on CUHK-PEDES, ICFG-PEDES and RSTPReid datasets, respectively. While there is a gap between our method and the current fully supervised SOTA method(RaSaBai et al. (2023)), our method is already very close to some fully supervised methods.
4.3 Ablation Study
In this chapter, we thoroughly analyze the effectiveness of each module in the CPCL method as shown in Table 4.
4.3.1 Prototypical Multi-modal Memory (PMM) and Hybrid-level Cross-modal Matching (HCM)
It should be noted that PMM and HCM need to be combined to work properly. HCM comprises two components: PCM and ICPM. As shown in Table 4, No.0 vs No.2, No.0 vs No.3, No.1 vs No.6, and No.1 vs No.5, we readily observe that the inclusion of the PMM and HCM proves beneficial for enhancing the model’s performance. Removing this module results in a certain degree of performance degradation.
4.3.2 Outlier Pseudo Label Mining (OPLM)
As shown in Table 4, No.0 vs No.1, No.2 vs No.6 and No.3 vs No.5, we conducted an ablation study of OPLM. OPLM makes significant contributions to the model’s performance. Being a plug-and-play component, OPLM can seamlessly integrate into weakly supervised TPRe-ID methods which are based on clustering algorithms. The removal of this module results in the inability to extract valuable diverse samples, leading to performance degradation of the model. More analysis of OPLM is shown in the Appendix.
4.3.3 cross-modadl PCM vs single-modal PCM
We attempt two types of Prototypical Contrastive Matching (PCM) to enforce the embedding of a sample to be more similar to its corresponding prototypes compared to other prototypes: cross-modal PCM and single-modal PCM, denoted as $PCM_{c}$ and $PCM_{s}$. The experimental results are shown No.0 vs No.3 in Table 5. When employing $PCM_{s}$, it achieves competitive performance on CUHK-PEDES. $PCM_{c}$ surpasses $PCM_{s}$ in terms of all metrics and we argue that $PCM_{c}$ promotes alignment between visual and textual samples. In conclusion, experiments prove that $PCM_{c}$ is more suitable for model learning.
4.3.4 Exploration of Momentum updating scheme for PMM
We employ a momentum updating scheme for PMM to update the prototypes and conduct experiments to highlight the necessity of a momentum updating scheme for PMM and its impact on the performance of CPCL, as illustrated in No.1 vs No.3 in Table 5. Disabling the momentum updating scheme eliminates prototypical feature variance and instability in the training process caused by update issues. However, it also hinders the model’s further performance. We argue that the momentum update of prototypical features enhances our method’s performance and adaptability. This is achieved by enabling the model to effectively capture changes and evolution in the data while preserving overall stability.
Figure 4: Visualization of Saliency Map based on Grad-CAM. The samples are from CUHK-PEDES.
Figure 5: t-SNE visualization of 10 random person samples in CUHK-PEDES. Text and image samples are denoted as triangles and circles, respectively. Samples with the same color indicate that they belong to the same person. The same letter means that it corresponds to the same person.
Figure 6: Comparison of the top-10 retrieved results on the CUHK-PEDES using the baseline (first row) and CPCL method (second row). The images corresponding to the query text, as well as the matched and mismatched images, are distinguished by green and red rectangles, respectively.
4.3.5 Two-stage OPLM vs One-stage OPLM
Although we propose a One-stage OPLM that can mine valuable unclustered samples, some image-text pairs still are unminable. Based on this fact, we propose a two-stage OPLM to make full use of these image-text pairs and conduct experiments to evaluate the effectiveness of the two-stage training scheme. As shown No.3 vs No.2 in Table 5, CPCL achieves better performance with the Two-stage OPLM training scheme. This phenomenon is intuitive, as the un-clustered image-text pairs are valuable and more training data allows the model to perform better on the test set.
4.4 Analysis of Qualitative Results
4.4.1 Visualization of Saliency Map
To better understand visual and textual alignment, we use Grad-CAMSelvaraju et al. (2017) to compute the saliency map based on the last attention layer’s [CLS] token of the image encoder. The objective is the cosine similarity between the encoding of the input image and the text encoding. As shown in Fig. 4, our method has guided the model to learn the semantic relevance of visual the textual modalities. Compared with the baseline, CPCL could focus more on attributes of human bodies described by the text.
4.4.2 Visualization of t-SNE
We use t-SNEVan der Maaten and Hinton (2008) to visualize the embeddings of image and text samples belonging to 10 random persons in CUHK-PEDES. As shown in Fig. 5, our approach has better intra-class compactness and inter-class separability the the baseline. The compactness and separability contribute to better weakly supervised TPRe-ID performance. We also can find that text samples are more diverse compared to image samples. This phenomenon is consistent with the reality of the data, as annotators describe the same image differently. This finding may indicate that the difficulty of the weakly supervised TPRe-ID lies in how to further process the text samples.
4.4.3 Visualization of retrieval results
We showcase the top-10 retrieval results obtained using textual descriptions from both the baseline and CPCL. As shown in Fig. 6, CPCL achieves more accurate retrieval results for a textual query. However, it is worth noting that our model also returned an incorrect retrieval result. This occurrence may be attributed to the global matching method employed in our model, which directly aligns global visual and textual embeddings and might overlook some crucial details. In future work, we are planning to alleviate this issue.
5 Conclusion
CPCL has shown considerable improvements on weakly supervised TPRe-ID. The core of the CPCL is to bring instance-level samples closer to their prototypes. Though the results are positive, there still exists a large gap from the fully supervised TPRe-ID, suggesting that the label by the clustering algorithm is not satisfactory enough. We argue that conducting pseudo labels refinery can further improve the performance of the model and we will further investigate it in our future work. Moreover, we hope that our proposed CPCL could inspire and facilitate follow-up works with potential.